# Take History as a Mirror in Heterogeneous Federated Learning

### v1

@@ Introduction
where $r$ represents the current communication round and $\mu\in(0,\frac{1}{r})$ is a constant. Eq. (15) modifies the $\ell_{2}$-norm of the aggregated result to match the average $\ell_{2}$-norm of all local gradients before aggregation, and it decreases as the training progresses. This adjustment empowers $\mathbf{INA}$ to prevent significant reductions in the $\ell_{2}$-norm of the aggregated result. Generally, a larger $\ell_{2}$-norm of gradient implies the presence of more knowledge, and the $\ell_{2}$-norm of gradient learned from the previously trained samples tend to be relatively smaller (Shin et al. 2022). Thus, in addition to smoothing the convergence process, $\mathbf{INA}$ also possesses the capability of perceiving the significance of updates.
Pipeline of FedHist
Algorithm 1 summarizes the pipeline of the server-side of FedHist. Initially, the server receives and caches the first $K$ local gradients (Lines 4-5), followed by the application of $\mathbf{EGS}$ to enhance their stability (Lines 6-8). Subsequently, the server obtains the utility value $\overline{\boldsymbol{Util}}$ through $\mathbf{HAA}$ and aggregates the $K$ local gradients, considering their staleness (Lines 13-14), followed by $\ell_{2}$-norm adjustment performed by $\mathbf{INA}$ (Line 15). After the server finishes updating the global model (Line 16) and distributes it to the participants (Line 17), it computes the utility value $\overline{\boldsymbol{Util}}$ of the clients participating in round $r-h$ (Lines 18-19) and then maintains a fixed length of the buffer (Lines 20-21).
@@
Table 1: Test Accuracy (%) when using FedHist and baselines on FMNIST, CIFAR-10, and CIFAR-100 Datasets. $N/K$ is considered at two different levels: 10 and 100, to assess different degrees of staleness. $\beta$ represents the degree of data heterogeneity, taking values of 0.3, 1.0, and $\infty$. The best result in each column is shown in bold and the second-best is underlined.
Figure 4: Performance of FedHist with different missing components on CIFAR-10.
Figure 5: Performance of FedHist and baselines on FMNIST with specific labels occurring only on a few slow clients.
Experiments
Experimental Settings
Datasets and models.
@@
Our experiments are conducted on 3 well-known datasets, including CIFAR-10, CIFAR-100 (Krizhevsky, Hinton et al. 2009), and Fashion MNIST (FMNIST) (Xiao, Rasul, and Vollgraf 2017). To better simulate real-world scenarios of IID and Non-IID data, we employ the Dirichlet distribution $Dir_{N}(\beta)$ (Hsu, Qi, and Brown 2019) to divide the dataset. Here, $\beta$ controls the degree of data heterogeneity, where a smaller $\beta$ signifies a more pronounced imbalance in the quantity and labels of samples among clients. Besides, we introduce different levels of staleness by setting $N/K$ (the ratio of the number of total clients to the participants in each round). This choice stems from a theorem stating that the minimum average staleness is approximately equal to $N/(2K)$ (Zhou et al. 2022). And all experiments are conducted using LeNet (LeCun et al. 1998).
Baselines.
We select several relevant algorithms for comparison: FedAvg (McMahan et al. 2017), DynSGD (Jiang et al. 2017), TWAFL (Chen, Sun, and Jin 2019), FedProx (Li et al. 2020), MOON (Li, He, and Song 2021), WKAFL (Zhou et al. 2022) and REFL (Abdelmoniem et al. 2023). While some of them were initially proposed for synchronous FL, we also acknowledge their potential to support asynchronous scenarios through simple modifications. Please refer to Appendix A, B for more details of settings and results.
Performance Comparison
Data heterogeneity.
@@
Table 2 and Figure 3 illustrate that FedHist exhibits faster convergence rates compared to the baselines across three levels of heterogeneity, while also demonstrating the highest test accuracy, especially in highly heterogeneous environments. This is mainly attributed to the $\mathbf{EGS}$ within FedHist, which enhances the stability of local gradients and mitigates the impact of Non-IID data. Moreover, it is evident that the test accuracy of all methods improves with the increase of $\beta$, highlighting the detrimental impact of data heterogeneity on the training process.
Degrees of staleness.
@@
Table 2 demonstrates the leading test accuracy of FedHist across various staleness levels. Although the test accuracy of all methods experiences a certain decline with the increase of $N/K$, the incorporation of the $\mathbf{HAA}$ in FedHist comprehensively considers the utility of client updates, thereby enabling its adaptability in highly stale environments and satisfactory robustness to staleness. Furthermore, WKAFL, REFL, and TWAFL achieve relatively commendable results owing to their concerted efforts in coping with staleness, while FedAvg performs the least favorably due to its lack of measures to address staleness.
Convergence speed.
@@
Table 2 demonstrates that to achieve the target accuracy (40%), FedHist achieves an acceleration ratio of approximately 1.8$\times$-2.0$\times$ compared to FedAvg, surpassing all baselines. The convergence trends of different methods displayed in Figure 3 intuitively highlight the advantages of FedHist in diverse data heterogeneity settings. More importantly, FedHist consistently gains a significant advantage in the early stages of training, and its convergence curve remains smoother. This is achieved by adaptively amplifying the $\ell_{2}$-norms of aggregated gradients through $\mathbf{INA}$.
@@
Table 2: Communication rounds in different methods to achieve target accuracy (40%) on CIFAR-10. R# represents the required number of communication rounds and S $\uparrow$ is the corresponding convergence speedup relative to FedAvg.
Figure 6: Performance of FedHist with different $h$.
Model Analysis
Ablation studies.
@@
To further investigate the effectiveness of each component of FedHist, we conduct ablation studies, as shown in Figure 4. We systematically remove each of the three core components of FedHist and denote it as w/o (without) to indicate the absence of that particular component. The results demonstrate that regardless of the various data heterogeneities or different model staleness scenarios, FedHist consistently exhibits rapid convergence capability and stability. Indeed, the inferior stability observed in FedHist (w/o $\mathbf{EGS}$) and FedHist (w/o $\mathbf{HAA}$) indicates the impact of Non-IID data and staleness leads to the deviation of local gradients, hinders the convergence of the global model. Also, the slower convergence speed observed in FedHist (w/o $\mathbf{INA}$) reflects the effectiveness of the $\ell_{2}$-norm amplification strategy. More details please refer to Appendix C.
Sensitivity of the buffer length.
We further analyze the impact of the buffer length $h$ on FedHist concerning different degrees of statistical heterogeneity. As shown in Figure 6, we plot the average test accuracy of the global model with standard deviation by searching $h$in {0, 1, 5, 10, 15, 20, 30, 50, 100}. Figure 6 shows that a larger $h$ results in performance degradation due to the significantly reduced value of utilizing gradients cached in the buffer for an extended period, thus causing the collaborative gradient selected by $\mathbf{EGS}$ to compromise the stability of local gradients, and a smaller $h$ would lead to reduced predictive reliability of $\mathbf{HAA}$. Especially, the inferior performance when $h$ = 1 highlights the superiority of FedHist over momentum-based approaches (Hsu, Qi, and Brown 2019; Li et al. 2019).
Impacts of the boosting factor.
@@
To validate the capability of FedHist in handling extremely heterogeneous scenarios, we conduct an experiment by setting the tenth class samples (for a 10-class classification task) to occur only in a small fraction of slow clients with an update frequency of 50% compared to normal clients, and compare the experimental results of all the methods. As shown in Figure 5, FedHist demonstrates the highest test accuracy for these samples, which remains unaffected by changes in different degrees of data heterogeneity or staleness. This can be attributed to its consideration of aggregation weights beyond staleness. Through this multi-dimensional weighting approach, FedHist is capable of adapting to extremely heterogeneous scenarios. Please refer to Appendix D for more details.
Conclusion
@@
In this work, we propose federated historical learning (FedHist), a novel asynchronous FL method that utilizes the server-side buffer to replay the knowledge of historical gradients and mirror the utility of clients. FedHist first enhances the stability of local gradients by fine-grained cross-aggregation with historical global gradients. Moreover, it assigns the aggregation weights in a multi-dimensional manner based on the utility and staleness of the clients and enables fast and stable convergence by adjusting the $\ell_{2}$-norm of the aggregated gradient. We conduct extensive experiments on public datasets and validate the effectiveness of FedHist in addressing the challenges posed by both Non-IID data and gradient staleness, simultaneously.

### v2

@@ II-D Intelligent $\ell_{2}$-Norm Amplification (INA)
By modifying the $\ell_{2}$-norm of the aggregated result to match the average of all local gradients’ before aggregation. This empowers $\mathbf{INA}$ to prevent reductions in the $\ell_{2}$-norm of the aggregated result. Generally, a larger $\ell_{2}$-norm of gradient implies the presence of more knowledge, and the $\ell_{2}$-norm of gradient learned from the previously trained samples tend to be relatively smaller [21]. Thus, in addition to smoothing the convergence process, $\mathbf{INA}$ also possesses the capability of perceiving the significance of updates.
TABLE I: Test Accuracy (%) when using FedHist and baselines on FMNIST, CIFAR-10, and CIFAR-100 Datasets. The best result in each column is shown in bold and the second-best is underlined.
@@
III Experiments
III-A Experimental Settings
@@
Our experiments are conducted on three well-known datasets, including CIFAR-10, CIFAR-100 [22], and Fashion MNIST (FMNIST) [23]. To better simulate real-world scenarios of IID and Non-IID data, we employ the Dirichlet distribution $Dir_{N}(\beta)$ [24] to divide the dataset. Here, $\beta$ controls the degree of data heterogeneity, where a smaller $\beta$ signifies a more pronounced imbalance in the quantity and labels of samples. And we introduce different levels of staleness by setting $N/K$. We select several relevant algorithms for comparison: FedAvg [1], DynSGD [25], TWAFL [17], FedProx [14], MOON [13], WKAFL [9] and REFL [26]. All experiments are conducted using LeNet [27].
III-B Performance Comparison
Data heterogeneity
@@ III-B Performance Comparison
Table I and Figure 3 illustrate that FedHist exhibits faster convergence rates compared to the baselines across three levels of heterogeneity, while also demonstrating the highest test accuracy, especially in highly heterogeneous environments. This is mainly attributed to the $\mathbf{EGS}$ within FedHist, which enhances the stability of local gradients and mitigates the impact of Non-IID data. Moreover, it is evident that the test accuracy of all methods improves with the increase of $\beta$, highlighting the detrimental impact of data heterogeneity on the training process.
Degrees of staleness
@@
Table I demonstrates the leading test accuracy of FedHist across various staleness levels. Although the test accuracy of all methods experiences a certain decline with the increase of $N/K$, the incorporation of the $\mathbf{HAA}$ in FedHist comprehensively considers the utility of client updates, thereby enabling its adaptability in highly stale environments and satisfactory robustness to staleness. Furthermore, WKAFL, REFL, and TWAFL achieve relatively commendable results owing to their concerted efforts in coping with staleness, while FedAvg performs the least favorably due to its lack of measures to address staleness.
Convergence speed
@@
Table II demonstrates that to achieve the target accuracy (40%), FedHist achieves an acceleration ratio of approximately 1.8$\times$-2.0$\times$ compared to FedAvg, surpassing all baselines. The convergence trends of different methods displayed in Figure 3 intuitively highlight the advantages of FedHist in diverse data heterogeneity settings. Also, FedHist consistently gains a significant advantage in the early stages of training, and its convergence curve remains smoother. This is achieved by adaptively amplifying the $\ell_{2}$-norms of aggregated gradients through $\mathbf{INA}$.
TABLE II: Communication rounds in different methods to achieve target accuracy (40%) on CIFAR-10. R# represents the required number of communication rounds and S $\uparrow$ is the corresponding convergence speedup relative to FedAvg.
@@
Figure 4: Performance with different missing components on CIFAR-10.
Figure 5: Performance with specific labels occurring only on a few slow clients.
III-C Model Analysis
Convergence analysis
Table II compares the convergence rate of FedHist with the aforementioned baselines. The results demonstrate in both IID and Non-IID scenarios, FedHist achieves the fastest convergence rate to reach the target accuracy. Similarly, REFL also achieves results that are second only to FedHist by employing similar techniques. In contrast, DynSGD, TWAFL, and WKAFL exhibit relatively poorer convergence rates due to their focus on selecting fresh gradients while overlooking the value of stale gradients. The convergence trends of different methods displayed in Figure 3 intuitively highlight the advantages of FedHist. From these convergence plots, it can be observed that as $\beta$ increases, the convergence speed improves for all methods.
Ablation studies
@@ III-C Model Analysis
To further investigate the effectiveness of each component of FedHist, we conduct ablation studies, as shown in Figure 4. We systematically remove each of the three components and denote it as w/o (without) to indicate the absence of that particular component. The results demonstrate that regardless of the various data heterogeneities or different model staleness scenarios, FedHist consistently exhibits rapid convergence capability and stability. Indeed, the inferior stability observed in FedHist (w/o $\mathbf{EGS}$) and FedHist (w/o $\mathbf{HAA}$) indicates the impact of Non-IID data and staleness leads to the deviation of local gradients, hinders the convergence of global model. The slower convergence speed observed in FedHist (w/o $\mathbf{INA}$) reflects the effectiveness of $\ell_{2}$-norm amplification strategy.
Verification of fairness improvement
@@
To validate the capability of FedHist in handling extremely heterogeneous scenarios, we conduct an experiment by setting the tenth class samples (for a 10-class classification task) to occur only in a small fraction of slow clients with an update frequency of 50% compared to normal clients, and compare the experimental results of all the methods. As shown in Figure 5, FedHist demonstrates the highest test accuracy for these samples, which remains unaffected by changes in different degrees of data heterogeneity or staleness. This can be attributed to its consideration of aggregation weights beyond staleness. Through this weighting approach, FedHist can adapt to heterogeneous scenarios while ensuring that slower clients also contribute, thereby promoting fairness.
IV Conclusion
@@
In this work, we propose a novel AFL method (FedHist) that replays the knowledge of historical gradients and mirrors the utility of clients. FedHist first enhances the stability of local gradients by fine-grained cross-aggregation. It assigns the aggregation weights in a multi-dimensional manner and enables fast and stable convergence by adjusting the $\ell_{2}$-norm of the aggregated gradient. We conduct extensive experiments on public datasets and validate its effectiveness.